An open economy is described by the following system of macroeconomic equations, in which all macroeconomic aggregate are measured in billions of Namibian dollars, N$:






 

(a)   Determine the equilibrium level of income/output.                                                                           (4)

1
Expert's answer
2021-05-16T17:24:32-0400

An open economy is described by the following system of macroeconomic equations, in which all macroeconomic aggregate are measured in billions of Namibian dollars, N$:

Y=C+I+G+X-M.

C=100+0.75Yd.

T=50+0.5y.

I=200.

X=200.

M=50+0.25y.

G=150.

(a)   Determine the equilibrium level of income/output


The calculation for the equilibrium level of output/income is as follows:

"C=100+0.75(Y-T)"

"=100+0.75[Y-(50+0.5Y)]"

"=100+0.75(Y-50-0.5Y)"

"=100+0.75Y-37.5-0.375Y"


"X-M=200-(50+0.25Y)"

"=200-50-0.25Y"

"=150-0.25Y"


Now put the values in the main equation of output.

"Y=C+I+G+X-M"

"Y=100+0.75Y-37.5-0.375Y+200+150+150-0.25Y"

"Y=100-37.5+200+150+150+0.75Y-0.375Y-).25Y"

"Y=562.5+0.125Y"

"Y=642.85"


Hence, the equilibrium level of income is $642.85.